Write the expression as a sum or difference of trigonometric functions:
2sin(35)cos(84)

Found 2 solutions by stanbon, Edwin McCravy:
Answer by stanbon(75887)

Write the expression as a sum or difference of trigonometric functions:
2sin(35)cos(84)
------
x+y = 84
x-y = 35
----
2x = 119
x = 59.5
y = 24.5
------
Ans:: 2sin(59.5-24.5)cos(59.5+24.5)
------
Cheers,
Stan H.
-------------------

Answer by Edwin McCravy(20059)

The other tutor doesn't understand. Here's what you want:
2sin(35°)cos(84°)

 sin(35°+84°) = sin(35°)cos(84°)+cos(35°)sin(84°)
 sin(35°-84°) = sin(35°)cos(84°)-cos(35°)sin(84°)

Simplify the left sides:

sin(35°+84°)= sin(119°) = sin(180°-61°) = sin(61°)
sin(35°-84°)= sin(-49°) = -sin(49°) 

         sin(61°) =  sin(35°)cos(84°)+cos(35°)sin(84°)
        -sin(49°) =  sin(35°)cos(84°)-cos(35°)sin(84°)
---------------------------------------------
sin(61°)-sin(49°) = 2sin(35°)cos(84°) 

The right side is what we were given:

Answer: sin(61°)-sin(49°)

Edwin